How many gallons of a 10% salt solution should be added to 20 gallons of a 60% salt solution to obtain a 50% salt solution?
Answer by addingup(3677)

0.10x+0.60(20) = 0.50(x+20)
0.10x+12 = 0.50x+10 substract 0.10x and 10 from both sides
2 = 0.40x flip it
0.40x = 2 divide both sides by 0.40
x = 5 gallons of the 60% solution should be added to the 20 gallon to obtain a 50% solution